# Blowing up Dirac’s theorem

### v1

@@ 3. Tools
For the proofs of Lemmata 4.1 and 2.2, we require a few basic results from probability and extremal hypergraph theory. An $s$-uniform hypergraph (or $s$-graph for short) is a pair $P=(V,E)$ where every $e\in E$ is an $s$-element subset of $V$. The minimum degree $\delta_{1}(P)$ of an $s$-graph $P=(V,E)$ is the minimum of $\deg_{P}(v)=|\{e\in E\colon v\in e\}|$ over all $v\in V$.
@@
A well-known probabilistic argument shows that relative minimum degree conditions are preserved with high probability when taking random induced subgraphs. Given a Dirac-type graph $G$, we use an $s$-uniform hypergraph $P$, called the property graph, which tracks the induced subgraphs that the minimum degree condition.
@@
The proof of Lemma 3.1 follows from a concentration analysis. For sake of completeness, the details are presented in section A.
@@
Our approach relies on the existence of complete blow-ups with clusters of polylogarithmic size in dense hypergraphs, as guaranteed by a classic result of Erdős [4]. We use a refinement of this result for graphs by Nikiforov [8, 9], which allows us to find blow-ups with clusters of logarithmic size, not just polylogarithmic. The following is an immediate corollary of [8, Theorem 1].
@@ 4. Simple blow-up covers
The proof of Lemma 2.1 is adapted from our work on hypergraphs [7, §7]. We obtain better bounds on the cluster sizes by applying Lemma 3.2 instead of the related result for hypergraphs by Erdős [4]. Let us start with the following result, which has appeared (with constant cluster size) in the setting of perfect tilings [6, Corollary 10.2].
@@
The next result allows us to pick up leftover vertices.
@@
Let $\varepsilon\gg 1/s\gg c\gg n$, and let $n^{1/2}\leq m_{1}\leq(\varepsilon/4)n$ and $m_{2}=c\log m_{1}$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Let $V\subseteq V(G)$ be an $m_{1}$-set. Then $G$ contains an $m_{2}$-balanced blow-up $T(\mathcal{W})$ such that $T$ has order $s$ and $\delta(T)\geq(1/2+\varepsilon/2)s$. Moreover, there is $W\in\mathcal{W}$ with $W\subseteq V$ and $W^{\prime}\cap V=\varnothing$ for all $W^{\prime}\in\mathcal{W}$ distinct from $W$.
We defer proofs of Lemmata 4.1 and 4.2 to section 5 and section 6 and continue to show the main result of this section.
@@
Introduce $c_{1},c_{2},c_{3}$ with $c_{1},\eta\gg c_{2}\gg c_{3}\gg 1/n$. Let $m_{i}=c_{i}\log n$, for $i\in\{1,2,3\}$. By Lemma 4.1, all but at most $\eta n$ vertices of $G$ may be covered with pairwise vertex-disjoint $m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{p}(\mathcal{V}_{p})\subseteq G$, where each $R_{i}$ has order $s$ and $\delta(G_{i})\geq(1/2+\varepsilon/2)s$. The uncovered vertices, denoted by $W$, are added onto blow-ups in two steps.
We first cover most vertices of $W$ as follows. While $|W|\geq\eta n/m_{1}$, iteratively apply Lemma 4.2 (with $|W|$ playing the rôle of $m_{1}$) to find vertex-disjoint blow-ups $T^{1}(\mathcal{W}_{1}),\dots,T^{p}(\mathcal{W}_{q})$ such that each $\mathcal{W}_{i}$ is an $m_{2}$-balanced set family with exactly one cluster contained in $W$. Each blow-up covers $(s-1)m_{2}$ vertices of $G-W$ and $m_{2}$ vertices of $W$. So the process requires at most $(s-1)|W|/m_{2}\leq n/m_{2}$ vertices of $G-W$, which means that we can greedily choose these blow-ups to be pairwise vertex-disjoint. Moreover, since $(s-1)m_{2}\leq\eta m_{1}$, we can also ensure that no cluster of $\mathcal{V}_{1}\cup\dots\cup\mathcal{V}_{p}$ is intersected in more than $(\eta/8)m_{1}$ vertices. Denote the remaining uncovered vertices by $U\subseteq W$.
To finish the covering, we apply Lemma 4.2 again to add each $u\in U$ onto a quasi $m_{3}$-balanced $\operatorname{\mathsf{P}}$-blow-up $S^{u}(\mathcal{U}_{u})$, where $S^{u}$ has order $s$ and $\{u\}\in\mathcal{U}_{u}$. (The requirements of the lemma can be met by replacing $u$ with $\varepsilon n$ copies of itself, which are dropped afterwards.) Note that this blow-up covers $(s-1)m_{3}$ vertices of $G-U$ and $1$ vertex of $U$. Since $(s-1)m_{3}|U|\leq\eta n$, we can select the blow-ups $S^{u}(\mathcal{U}_{u})$ ($u\in W$) to be pairwise vertex-disjoint. Moreover, since $(s-1)m_{3}\leq\eta m_{2}$, we may ensure that any cluster of $\mathcal{V}_{1},\dots,\mathcal{V}_{p},\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{q}$ is intersected in at most $(\eta/8)m_{2}$ vertices.
@@
We delete the vertices of each $\mathcal{W}_{i}$ from each $\mathcal{V}_{j}$, keeping the names for convenience. Afterwards, we delete the vertices from $\mathcal{U}_{\ell}$ from each of $\mathcal{W}_{i}$ and $\mathcal{V}_{j}$, keeping again the names. To finish, we need to do two things: since the clusters of the families $\mathcal{W}_{i}$ and $\mathcal{V}_{i}$ are still much larger compared to $\mathcal{U}_{i}$, we need to split each of those clusters to make them smaller. Secondly, to obtain quasi balanced blow-ups, we need to create singleton clusters in those blow-ups. We address both of those issues as follows. First, we split each $\mathcal{W}_{i}$ and each $\mathcal{V}_{j}$ into $(1\pm\eta/2)((s-1)m_{3})$-balanced families. (This can be done greedily.) then, we partition each of these $(1\pm\eta/2)((s-1)m_{3})$-balanced families into $s$ quasi $(1\pm\eta)m_{3}$-balanced families. To do this, assuming the family is $\{X_{1},\dotsc,X_{s}\}$, can partition each $X_{i}$ into $s$ sets, one of size $1$, and the other ones of size as equal as possible, that is, of size $(1\pm\eta)m_{3}$. Then we form $s$ new families by including in each one exactly one $1$-sized set, and $s-1$ many $(1\pm\eta)m_{3}$-sized sets. Hence every family is now quasi $(1\pm\eta)m_{3}$-balanced, and we can finish with $c=c_{3}$. ∎
@@ 5. Almost blow-up covers
Here we show Lemma 4.1. We require the following notion of (weak) quasirandomness for $s$-uniform hypergraphs. For an $s$-graph $P$, the density of a tuple $(V_{1},\dots,V_{k})$ of pairwise disjoint vertex sets is
@@
We say that $(V_{1},\dots,V_{k})$ is $(\varrho,d)$-lower-regular if for $d=e_{P}(V_{1},\dots,V_{k})$ and all choices of $X_{1}\subseteq V_{1},$ $\dots$,… , $X_{k}\subseteq V_{k}$ satisfying $|X_{1}|\geq\varrho|V_{1}|,\,\dots,\,|X_{k}|\geq\varrho|V_{k}|$, we have $d_{P}(X_{1},\dots,X_{k})\geq d-\varrho$. Denote by $\mathcal{Q}(s,m,\varrho,d)$ the set of $(\varrho,d^{\prime})$-lower-regular $s$-tuples which are $m^{\prime}$-balanced, with $m^{\prime}\geq m$, and $d^{\prime}\geq d$.
Our next lemma shows that one can partition most of the vertices of a hypergraph with large enough minimum degree into balanced lower-regular tuples. This is a straightforward consequence of the Weak Hypergraph Regularity Lemma [6, Theorem B.2]. A simpler, alternative proof of this fact appears in our related work on hypergraphs [7, §7.3]. For sake of completeness, the details are presented in section B.
@@
By Lemma 5.1, the $s$-graph $P$ contains a $\mathcal{Q}(s,\alpha n,\varrho,\eta)$-tiling that covers all but $\eta n$ vertices. Consider one of the $(\varrho,\eta)$-lower-regular $s$-tuples with $\alpha n$-balanced parts $\mathcal{V}=\{V_{1},\dots,V_{s}\}$. Note that each $\mathcal{V}$-partite edge of $P$ corresponds to some labelled $s$-vertex graph with minimum degree at least $(1/2+\varepsilon/2)s$. Note that there are at most $2^{s^{2}}$ different labelled $s$-vertex graphs. By the pigeonhole principle there is an $s$-vertex graph $R_{1}$ with $\delta(R_{1})\geq(1/2+\varepsilon/2)s$ such that at least $(\eta-\varrho)2^{-s^{2}}(\alpha n)^{s}$ edges of $P$ correspond to a labelled copy of $R_{1}$. By Lemma 3.2, we may find a subgraph $R_{1}(m)\subseteq G$ with $\mathcal{V}$-partite edges, which constitutes one of the desired blow-ups. We repeat this procedure as long as there are at least $(\eta-\varrho)(\varrho\alpha n)^{s}$ $\mathcal{V}$-partite unused edges, thus obtaining pairwise vertex-disjoint blow-ups $R_{1}(m),\dots,R_{\ell}(m)$. Since the tuple $(V_{1},\dots,V_{s})$ is $(\varrho,\eta)$-lower-regular, this procedure stops when all but at most $\varrho\alpha n$ vertices of $\mathcal{V}$ are used in each cluster of $\mathcal{V}$. After iterating this over all regular tuples, we have thus covered all but $\varrho n+\eta n\leq 2\eta n$ vertices, which allows us to conclude with $2\eta$ in place of $\eta$. ∎

### v2

@@ 3. Tools
For the proofs of Lemmata 4.1 and 2.2, we require a few basic results from probability and extremal hypergraph theory. An $s$-uniform hypergraph (or $s$-graph for short) $P$ consists of vertices $V(P)$ and edges $E(P)$, where every edge $e\in E(P)$ is a set of $s$ vertices. The minimum degree $\delta_{1}(P)$ is the largest $m$ such that every vertex of $P$ is on at least $m$ edges.
@@
It is well-known that sampling an induced subgraph in a Dirac-type graph approximately preserves minimum degree conditions. We use an $s$-graph $P$, called the property graph, to track this phenomenon locally.
@@
The proof of Lemma 3.1 follows from a concentration analysis. For the sake of completeness, the details are presented in section A.
@@
Our approach relies on the existence of complete blow-ups with clusters of polylogarithmic size in dense hypergraphs, as guaranteed by a classic result of Erdős [5]. We use a refinement of this result for graphs by Nikiforov [12, 13], which allows us to find blow-ups with clusters of logarithmic size, not just polylogarithmic. The following is an immediate corollary of [12, Theorem 1].
@@ 4. Simple blow-up covers
In this section, we show Lemma 2.1. The proof relies on ideas from our work on hypergraphs [11, §7], where a similar lemma is proved. Here we obtain better bounds on the cluster sizes by applying Lemma 3.2 instead of the related result for hypergraphs by Erdős [5]. We derive Lemma 2.1 from the following two results. The first one allows us to cover all but a linear number of vertices with blow-ups. It first appeared with constant cluster size in the setting of perfect tilings [10, Corollary 10.2].
@@
The next result allows us to pick up the remaining vertices.
@@
Let $\varepsilon\gg 1/s\gg c\gg n$, and let $n^{1/2}\leq m_{1}\leq(\varepsilon/4)n$ and $m_{2}=c\log m_{1}$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Let $X\subseteq V(G)$ be an $m_{1}$-set. Then $G$ contains an $m_{2}$-balanced blow-up $T(\mathcal{W})$ such that $T$ has order $s$ and $\delta(T)\geq(1/2+\varepsilon/2)s$. Moreover, there is $W\in\mathcal{W}$ with $W\subseteq X$ and $W^{\prime}\cap X=\varnothing$ for all $W^{\prime}\in\mathcal{W}$ distinct from $W$.
We defer proofs of Lemmata 4.1 and 4.2 to section 5 and section 6 and continue to show Lemma 2.1. To explain the approach, consider constants $c_{1}\gg c_{2}\gg c_{3}$, and let $G$ be a graph as in the statement. We work with clusters of sizes $m_{i}=c_{i}\log n$. To begin, take $m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{p}(\mathcal{V}_{p})\subseteq G$ obtained from Lemma 4.1 covering all but $o(n)$ vertices. In particular, each $R_{i}$ has $s$ vertices and inherits the minimum degree condition up to a small error. It remains to include the uncovered vertices, which we denote by $X$. To this end, we iteratively apply Lemma 4.2 to find additional disjoint $m_{2}$-balanced blow-ups $T_{1}(\mathcal{W}_{1}),\dots,T_{p}(\mathcal{W}_{q})$ each covering $m_{2}$ vertices of $X$ that have (in total) only a few vertices in each cluster of the families $\mathcal{V}_{i}$. As before, each $T_{j}$ has $s$ vertices and approximately inherits the minimum degree condition. We delete the vertices of each $\mathcal{W}_{j}$ from each $\mathcal{V}_{i}$ to ensure that the families are disjoint. Note that this leaves the families $\mathcal{V}_{i}$ still nearly balanced. At the end of this process, only $O(n^{3/4})$ vertices of $Y\subseteq X$ remain uncovered. To finish, we cover each $u\in Y$ with a singleton cluster of a quasi $m_{3}$-balanced blow-up $S_{u}(\mathcal{U}_{u})$, which is again possible by an application of Lemma 4.2. Here, each $S_{\ell}$ has $s+1$ vertices and approximately inherits the minimum degree condition. As before, one can ensure that the families $\mathcal{U}_{u}$ are disjoint from the families $\mathcal{V}_{i}$ and $\mathcal{W}_{j}$ by controlling the intersection and increasing the imbalance of $\mathcal{V}_{i}$ and $\mathcal{W}_{j}$ slightly. This leaves us with a complete cover of $G$ using three types of families, whose clusters are of size roughly $m_{1}$, roughly $m_{2}$ and, lastly, of size $m_{3}$ with one singleton cluster. To finish, we subdivide the larger families until all families have one singleton cluster and $s$ other clusters, which are all of size roughly $m_{3}$. We remark that only the third family we obtain is quasi balanced and not balanced, and therefore adding the singleton cluster to the sets in the first two families is somewhat unnecessary. However, doing so helps us to simplify the outcome, because we obtain a partition made only of the same type of blow-ups. Let us now turn to the details.
@@
Introduce $c_{1},c_{2},c_{3}$ with $\eta\gg c_{1}\gg c_{2}\gg c_{3}\gg 1/n$. Let $m_{i}=c_{i}\log n$, for $i\in\{1,2,3\}$. By Lemma 4.1 (with $\eta^{3}$ playing the rôle of $\eta$), all but at most $\eta^{3}n$ vertices of $G$ may be covered with pairwise vertex-disjoint $m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{p}(\mathcal{V}_{p})\subseteq\penalty 10000\ G$, where each $R_{i}$ has order $s$ and $\delta(R_{i})\geq(1/2+\varepsilon/2)s$. The uncovered vertices, denoted by $X$, are added onto blow-ups in two steps. We first include most vertices of $X$ as follows.
Claim 4.3.
For $q\leq\eta^{3}n/m_{2}$, there are vertex-disjoint blow-ups $T_{1}(\mathcal{W}_{1}),\dots,T_{q}(\mathcal{W}_{q})$ such that
[…]
    each $\mathcal{W}_{i}$ is an $m_{2}$-balanced set family with at least one cluster contained in $X$,
[…]
    each $T_{i}$ is an $s$-vertex graph with $\delta(T_{i})\geq(1/2+\varepsilon/2)s$,
[…]
    $\bigcup_{i}\bigcup\mathcal{W}_{i}$ covers at most $(\eta/4)m_{1}$ vertices of any cluster of $\mathcal{V}_{1},\dots,\mathcal{V}_{p}$ and
[…]
    $\bigcup_{i}\bigcup\mathcal{W}_{i}$ covers all but at most $n^{3/4}$ vertices of $X$.
Proof of the claim.
We construct the blow-ups one after another, covering at least $m_{2}$ vertices of $X$ in each step until all but at most $n^{3/4}$ vertices are covered. This process stops after at most $q=\eta^{3}n/m_{2}$ steps. Suppose we have already constructed blow-ups $T_{1}(\mathcal{W}_{1}),\dots,T_{\ell}(\mathcal{W}_{\ell})$ for $1\leq\ell<q$ satisfying part ii and iii, which cover the vertex set $W=\bigcup_{j=1}^{\ell}\bigcup\mathcal{W}_{j}$ and at least $\ell m_{2}$ vertices of $X$. If $|X\smallsetminus W|\leq n^{3/4}$, we stop the process and finish with $\ell$ playing the rôle of $q$. Otherwise, we shall find a blow-up $T_{\ell+1}(\mathcal{W}_{\ell+1})\subseteq G-W$ satisfying part ii and iii, which in addition covers $m_{2}$ vertices of $X$. Indeed, let us call a cluster $V$ of a set family $\mathcal{V}_{i}$ saturated if $V$ intersects in at least $(\eta/8)m_{1}$ elements with the vertices of $W$. Note that $|W|\leq sm_{2}q\leq s\eta^{3}n\leq\eta^{2}n\leq(\varepsilon/4)n$. So there are at most $h=\eta^{2}n/(\eta m_{1}/8)\leq 8\eta n/m_{1}$ saturated clusters. Denote by $B\subseteq V(G)$ the union of the vertices of the saturated clusters. It follows that $|B|\leq hm_{1}\leq 8\eta n\leq(\varepsilon/4)n$. Let $G^{\prime}=G-W-B$, and note that $G^{\prime}$ has minimum degree at least $(1/2+\varepsilon/2)n$. We may thus apply Lemma 4.2 with an $\lfloor n^{3/4}\rfloor$-set $X^{\prime}\subseteq X$ playing the rôle of $X$, and $2c_{2}$ playing the rôle of $c$. Since $2c_{2}\log(n^{3/4})\geq m_{2}$, this gives a blow-up $T_{\ell+1}(\mathcal{W}_{\ell+1})$ such that each $\mathcal{W}_{i}$ is an $m_{2}$-balanced set family with exactly one cluster contained in $X^{\prime}\subseteq X$. Clearly, $T_{\ell+1}(\mathcal{W}_{\ell+1})$ is disjoint from $W=\bigcup_{j=1}^{\ell}\bigcup\mathcal{W}_{j}$. Moreover, $T_{\ell+1}$ is an $n$-vertex graph satisfying $\delta(T_{i})\geq(1/2+\varepsilon/2)s$, which gives part ii. Let $W^{\prime}=W\cup\bigcup\mathcal{W}_{\ell+1}$. Then $W^{\prime}$ intersects saturated clusters in at most $(\eta/4)m_{1}$ vertices and unsaturated clusters in at most $(\eta/8)m_{1}+sm_{2}\leq(\eta/4)m_{1}$ vertices, in accordance with part iii. ∎
Fix $T_{1}(\mathcal{W}_{1}),\dots,T_{q}(\mathcal{W}_{q})$ as in Claim 4.3 and denote by $Y\subseteq X$ the set of at most $n^{3/4}$ uncovered vertices. Denote by $W$ the set of vertices covered by $\mathcal{W}_{1},\dots,\mathcal{W}_{q}$, and observe that $|W|\leq qsm_{2}\leq\eta^{2}n$. We finish the covering at follows.
Claim 4.4.
There is a family of pairwise disjoint blow-ups $S_{y}(\mathcal{U}_{y})$ for $y\in Y$ that together cover $Y$ such that each $S_{i}$ is a graph on $s+1$ vertices with $\delta(S_{i})\geq(1/2+\varepsilon/2)s$. Moreover, $U=\bigcup_{y\in Y}\bigcup\mathcal{U}_{y}$ is disjoint from $W$ and has at most $(\eta/4)m_{1}$ vertices of any cluster of $\mathcal{V}_{1},\dotsc,\mathcal{V}_{p}$.
Proof of the claim.
We construct the blow-ups one after another. Suppose we have so far covered vertices $Y^{\prime}\subseteq Y$ with blow-ups $\{S_{y}(\mathcal{U}_{y})\}_{y\in Y^{\prime}}$, where each blow-up $S_{y}(\mathcal{U}_{y})$ covers $y$ and no other vertex from $Y$. Suppose furthermore that $U^{\prime}=\bigcup_{y\in Y^{\prime}}\bigcup\mathcal{U}_{y}$ is disjoint from $W$ and $U^{\prime}$ has at most $(\eta/4)m_{1}$ vertices of any cluster of $\mathcal{V}_{1},\dotsc,\mathcal{V}_{p}$. We have $|U^{\prime}|\leq|Y|(s+1)m_{2}\leq(s+1)n^{3/4}\log n\leq\varepsilon n/16$. Say that a cluster $V$ of $\bigcup_{i}\mathcal{V}_{i}$ is saturated if $|V\cap U^{\prime}|\geq(\eta/8)m_{1}$. Denote by $h$ the number of saturated clusters and note that $h\leq|U^{\prime}|/(\eta m_{1}/8)$. Hence, the union $D\subseteq V(G)$ of the vertices of saturated clusters has size at most $hm_{1}\leq(\varepsilon/16)n$. So we have
$|W|+|U^{\prime}|+|Y|+|D|\leq\eta^{2}n+\varepsilon n/16+n^{3/4}+\varepsilon n/16\leq(\varepsilon/4)n\,.$
Fix an uncovered vertex $y\in Y\smallsetminus Y^{\prime}$. Let $G^{\prime}$ be an auxiliary graph obtained from $G-W-U^{\prime}-Y-D$ by replacing $y$ with a set of $n^{3/4}$ copies $X$ of itself. Note that $\delta_{1}(G^{\prime})\geq(1/2+\varepsilon/2)v(G^{\prime})$. We may thus apply Lemma 4.2 to find an $m_{3}$-balanced blow-up $S_{y}(\hat{\mathcal{U}}_{y})$, where $S$ has $s+1$ vertices and satisfies $\delta(S_{y})\geq(1/2+\varepsilon/2)s$. Delete from $\hat{\mathcal{U}}_{y}$ all but one copy of $y$ to obtain a quasi $m_{3}$-balanced family $\mathcal{U}_{y}$. Clearly, $S_{y}(\mathcal{U}_{y})$ covers exactly one vertex in $Y$ and is disjoint from the blow-ups $\{S_{y^{\prime}}(\mathcal{U}_{y^{\prime}})\}_{y^{\prime}\in Y^{\prime}}$ that were selected before. Moreover, the clusters $\mathcal{U}_{y}$ are disjoint from $W$ and have at most $sm_{3}+1\leq(\eta/8)m_{1}$ vertices in any unsaturated cluster of $\mathcal{V}_{1},\dots,\mathcal{V}_{p}$. Moreover, $\mathcal{U}_{y}$ is disjoint from the saturated clusters. It follows that $\bigcup\mathcal{U}_{y}\cup U^{\prime}$ has at most $(\eta/4)m_{1}$ vertices of any cluster of $\mathcal{V}_{1},\dotsc,\mathcal{V}_{p}$. ∎
Fix blow-ups $S_{y}(\mathcal{U}_{y})$ for $y\in Y$ as in Claim 4.4. Let $U=\bigcup_{y\in Y}\bigcup\mathcal{U}_{y}$. We delete the vertices of each $\mathcal{W}_{i}$ and $\mathcal{U}_{y}$ from each $\mathcal{V}_{j}$, keeping the names for convenience. Note that afterwards, $\mathcal{V}_{j}$ is still $(1\pm\eta/2)m_{1}$-balanced by Claim 4.3iii and Claim 4.4.
@@
To finish, we have to address the following two issues. Firstly, the clusters of the families $\mathcal{W}_{i}$ and $\mathcal{V}_{i}$ are still much larger compared to $\mathcal{U}_{i}$. Secondly, the families do not have singleton clusters as required for quasi balancedness. To deal with this, we first split each $\mathcal{W}_{i}$ and each $\mathcal{V}_{j}$ into $(1\pm\eta/2)((s-1)m_{3})$-balanced families. (This can be done greedily.) Then, we partition each of these $(1\pm\eta/2)((s-1)m_{3})$-balanced families into $s$ quasi $(1\pm\eta)m_{3}$-balanced families. Indeed, assuming the family is $\{X_{1},\dotsc,X_{s}\}$, one can partition each $X_{i}$ into $s$ sets, one of size $1$, and the other ones of size as equal as possible, that is, of size $(1\pm\eta)m_{3}$. Then we form $s$ new families by including in each one exactly one singleton and s1s-1 many $(1\pm\eta)m_{3}$-sets. Hence every family is now quasi $(1\pm\eta)m_{3}$-balanced, and we can finish with $c=c_{3}$. Lastly, note that each of the reduced graphs $R_{i}$, $T_{j}$ and $S_{y}$ still has minimum degree at least $(1/2+\varepsilon/4)(s+1)$ after this process. ∎
@@ 5. Almost blow-up covers
Here we show Lemma 4.1, which allows us to cover most vertices of a Dirac-type graph $G$ with blow-ups whose reduced graphs approximately inherit the minimum degree condition. Let us outline the proof. Consider the property $s$-graph $P$ defined in Lemma 3.1, which tracks the $s$-sets $S\subseteq V(G)$ that inherit the minimum degree conditions of $G$. By degree inheritance, $P$ has large enough minimum degree to contain an almost perfect matching. In fact, we shall prove in Lemma 5.1 that one can even cover most vertices of $P$ with pairwise disjoint quasirandom $s$-partite $s$-graphs. Afterwards, we cover most of the vertices of each quasirandom $s$-graph separately with suitable blow-ups using Lemma 3.2.
To formalise this discussion, we introduce the following notion of (weak) quasirandomness for $s$-uniform hypergraphs. For an $s$-graph $P$, the density of a tuple $(V_{1},\dots,V_{k})$ of pairwise disjoint vertex sets is
@@
We say that $(V_{1},\dots,V_{k})$ is $(\varrho,d)$-lower-regular if for $d=d_{P}(V_{1},\dots,V_{k})$ and all choices of $X_{1}\subseteq V_{1},$ $\dots,$ $X_{k}\subseteq V_{k}$ satisfying $|X_{1}|\geq\varrho|V_{1}|,\,\dots,\,|X_{k}|\geq\varrho|V_{k}|$, we have $d_{P}(X_{1},\dots,X_{k})\geq d-\varrho$. Denote by $\mathcal{Q}(s,m,\varrho,d)$ the set of $(\varrho,d^{\prime})$-lower-regular $s$-tuples which are $m^{\prime}$-balanced, with $m^{\prime}\geq m$, and $d^{\prime}\geq d$.
Our next lemma shows that one can partition most of the vertices of a hypergraph with large enough minimum degree into balanced lower-regular tuples. This is a straightforward consequence of the Weak Hypergraph Regularity Lemma [10, Theorem B.2]. A simpler, alternative proof of this fact appears in our related work on hypergraphs [11, §7.3]. For the sake of completeness, the details are presented in section B.
@@
By Lemma 5.1, the $s$-graph $P$ contains a $\mathcal{Q}(s,\alpha n,\varrho,\eta)$-tiling that covers all but $\eta n$ vertices. Consider one of the $(\varrho,\eta)$-lower-regular $s$-tuples with $\alpha n$-balanced parts $\mathcal{V}=\{V_{1},\dots,V_{s}\}$. Note that each $\mathcal{V}$-partite edge of $P$ corresponds to some labelled $s$-vertex graph with minimum degree at least $(1/2+\varepsilon/2)s$. Note that there are at most $2^{s^{2}}$ different labelled $s$-vertex graphs. By the pigeonhole principle and $(\varrho,\eta)$-lower-regularity, there is an $s$-vertex graph $R_{1}$ with $\delta(R_{1})\geq(1/2+\varepsilon/2)s$ such that at least $(\eta-\varrho)2^{-s^{2}}(\alpha n)^{s}$ edges of $P$ correspond to a labelled copy of $R_{1}$. By Lemma 3.2, we may find an $m$-balanced blow-up $R_{1}(\mathcal{V}_{1})\subseteq G$ with $\mathcal{V}$-partite edges, which constitutes one of the desired blow-ups. We repeat this procedure as long as there are at least $(\eta-\varrho)(\varrho\alpha n)^{s}$ $\mathcal{V}$-partite unused edges, thus obtaining pairwise vertex-disjoint blow-ups $R_{1}(m),\dots,R_{\ell}(m)$. Since the tuple $(V_{1},\dots,V_{s})$ is $(\varrho,\eta)$-lower-regular, this procedure stops when all but at most $\varrho\alpha n$ vertices of $\mathcal{V}$ are used in each cluster of $\mathcal{V}$. After iterating this over all regular tuples, we have thus covered all but $\varrho n+\eta n\leq 2\eta n$ vertices, which allows us to conclude with $2\eta$ in place of $\eta$. ∎